OpenStudy (anonymous):

find a equation for the line that contains the point (-2,3) and has a slope 5/7.

OpenStudy (anonymous):

you can use the point slope form. \[y - y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

Label your coordinate point (x1, y1) and your slope m, then substitute into the formula

OpenStudy (anonymous):

y=mx+b y=5/7 x+b 3=5/7 (-2)+b

OpenStudy (anonymous):

y=mx+c 3=5*(-2)/7+c c=31/7 y=mx+31/7

OpenStudy (anonymous):

or sub into the slope intercept form like the wiz diz

OpenStudy (anonymous):

solve for b 3=-10/7+b 3 +10/7=b 4 (3/7)=b

OpenStudy (anonymous):

y=5/7 x +4 3/7

OpenStudy (anonymous):

"Like the wiz diz." <3